the diagnol of a rectangle is eighteen inches more than its width. The length of the rectangle is seventeen more than its width.what are the demensions of the rectangle
Answer by checkley79(3341)

w^2+(w+17)^2=(w+18)^2
w^2+w^2+34w+289=w^2+36w+324
2w^2-w^2+34w-36w+289-324=0
w^2-2w-35=0
(w-7)(w+5)=0
w-7=0
w=7 ans. for the wwidth.
Proof:
7^2+(7+17)^2=(7+18)^2
49+24^2=25^2
49+576=625
625=625